# Rates of Convergence of Generalised Variational Inference Posteriors under Prior Misspecification

### v1

@@ 5.5 Rate of Convergence
$\frac{M}{\beta}n^{-1}\geq Q_{n}N_{\varepsilon_{n}}^{c}(L_{n}(N_{\varepsilon_{n}}^{c})-L_{n}(\Theta))\geq 0\,.$
@@
5.6 Extensions
We collect here the proofs of the results from Section 3.6.
5.6.1 Existence and Uniqueness
Since Lemma 2 applies regardless of the divergence, with some more work we arrive at the conclusion of Proposition 8. In proving existence, we follow the ideas of [29] where they prove existence of minimisers for Kullback–Leibler divergence minimisation to some probability measure over a subset of possible probability measures over infinite dimensional Polish spaces.
Proof of Proposition 8.
As the loss is lower bounded, we may let this constant be $c:=\inf_{\theta}L(X_{1}^{n},\,\theta)$. Suppose without loss of generality that this constant is negative. Then we claim that $\forall C>c$ the sublevel sets of the GVI functional $T_{n}(\cdot)$ are weak compact. As the loss and the divergence are lower semi–continuous, it follows that $\nu\mapsto J_{L_{n}}(\nu)$ is lower semi–continuous and so $\nu\mapsto T_{n}(\nu):=J_{L_{n}}(\nu)+D(\nu:\Pi)$ is lower semi–continuous [11]. Hence, by Prokhorov’s theorem we need to show that sublevel sets are tight. This is immediate if $\Theta$ is compact, so we now show the result for $\Theta$ that are not.
[…]
$\liminf_{k}T_{n}(\nu_{k})\geq T_{n}(\nu_{\star})$
For the uniqueness part, suppose that there exist two minimisers, $\nu_{1}$ and $\nu_{2}$. Then, as $\mathcal{Q}$ is convex we may pick $a\in(0,1)$ and let
$\nu:=a\nu_{1}+(1-a)\nu_{2}$ (13)
such that
$=aT_{n}(\nu_{1})+(1-a)T_{n}(\nu_{2})$
where the inequality follows by the convexity of the divergence. But as $\nu_{1}$ and $\nu_{2}$ are both minimisers $T_{n}(\nu_{1})=T_{n}(\nu_{2})$ so $\nu$ would further lower the GVI objective, so these cannot be minimisers. Hence, we conclude that the only way Equation 13 may hold is if $\nu_{1}=\nu_{2}$ (in the sense of measures), and the minimiser may not be written as a convex combination of elements of $\mathcal{Q}$ so this minimiser is an extreme point, and unique. ∎
5.6.2 Relaxation of $M$ and $\beta$
Proof of Proposition 9.
The result is immediate from the corresponding proofs of Theorems 6 and 7. ∎
5.6.3 Concentration of GVI Posterior Measures under Unbounded Divergences
@@
By Proposition 8, $Q_{n}$ exists under unbounded divergences. Then, for any such sequence $(\nu_{n})\subset\mathcal{Q}$ it follows by assumption that $P_{0}$ almost surely
$\frac{1}{n}T_{n}(\nu_{n})=J_{L_{n}}(\nu_{n})+(n\beta)^{-1}D(\nu_{n}:\Pi)\rightarrow L(\Theta)\,.$
$\frac{1}{n}T_{n}(\nu_{n})\geq\frac{1}{n}T_{n}(Q_{n})\geq J_{L_{n}}(Q_{n})=\int_{A}L_{n}(\theta)Q_{n}(d\theta)+\int_{A^{c}}L_{n}(\theta)Q_{n}(d\theta)\geq L_{n}(\Theta)\,.$
Note also that
$\int_{A}L_{n}(\theta)Q_{n}(d\theta)+\int_{A^{c}}L_{n}(\theta)Q_{n}(d\theta)\geq Q_{n}A\>L_{n}(A)+Q_{n}A^{c}\,L_{n}(A^{c})\geq L_{n}(\Theta)\,.$
By the lower semi–continuity of $L_{n}$ we have that $\exists\theta\in A^{c}$ for which $L(\theta)=L(\Theta)$ and therefore $L_{n}(A^{c})\rightarrow L(\Theta)$ $P_{0}$–a.s. Therefore,
$Q_{n}A\>L_{n}(A)+Q_{n}A^{c}\,L_{n}(A^{c})\geq Q_{n}A\>L_{n}(A)+Q_{n}A^{c}\,L_{n}(\Theta)\geq L_{n}(\Theta)$
Taking limits, it follows $P_{0}$ almost surely that $J_{L_{n}}(Q_{n})\rightarrow L(\Theta)$, and as $\liminf_{n}L_{n}(A)>L(\Theta)$ $P_{0}$–a.s. by assumption, it must follow that $Q_{n}A\rightarrow 0$ with $P_{0}$ probability 1. ∎
For unbounded divergences we will generally not have that $J_{L_{n}}(\nu_{n})=L_{n}(\Theta)$ at any finite $n$ as any atomic measure will have infinite divergence to any non–atomic prior. Therefore, we are unable to employ the same methodology in proving rates of convergence for GVI posterior measures under unbounded divergences as in the proof of Theorem 7.
6 Concluding Remarks
@@
It is worth noting that these results have further impact to Federated GVI [27], the study of GVI posteriors over a network of clients with private, local data sets. In particular such FedGVI posteriors will exist under the assumptions presented here through the different choices of loss functions and priors. Moreover, a global prior of some uniformed server, given a bounded divergence, may not skew the global posterior after incorporating local data towards unfavourable hypotheses that do not agree.
Finally, while we were able to derive existence and uniqueness of GVI posteriors under unbounded divergences, rates of convergence of GVI posterior measures under unbounded divergences remains an open challenge.
Appendix A Supplementary Definitions
@@
where the first is a Bochner integral (see [29]).
@@ Appendix B Classical Robustness to Prior Misspecification
We briefly discuss the relation of our work to robustness to prior misspecification in the sense of Bayesian sensitivity analysis [6, 17]. This line of work has examined the stability of Bayesian posteriors under small perturbations of the prior with respect to some other measure, for instance through additive Huber contamination as mentioned in Section 2.2. In particular, this has been used to select a robust prior for Bayes which leads to optimal decision making under potential perturbations of the prior measure. In this paper however, we aim to answer the reverse: under what conditions can GVI lead to optimal decision making under ill chosen priors.
Local sensitivity of posteriors generated with $\varepsilon$–contaminated priors, for instance by the additive Huber contamination model $\mathcal{G}=\{\Pi_{\varepsilon}:=(1-\varepsilon)\Pi_{0}+\varepsilon G:\varepsilon\in(0,1)\}$, is often studied [17] through the Gâteux differential
@@
where $d(\cdot:\cdot)$ is some distance or divergence measure; Gustafson and Wasserman [17] for instance use the total variation distance. Here $\Pi_{0}$ is the uncontaminated prior, $\Pi_{\varepsilon}$ the $\varepsilon$–contaminated prior in the direction of $G$, and $Q^{n}_{0}$ and $Q^{n}_{\varepsilon}$ the respective Bayes posteriors. The unrestricted choice of a divergence other than the Kullback–Leibler in GVI generally prohibits explicit expressions of $Q^{n}_{0}$ and $Q^{n}_{\varepsilon}$ that demonstrate the the exact dependence on the prior, which makes this approach impracticable for our purposes.
This presents an interesting theoretical question to the stability of Bayesian posteriors with respect to prior misspecification and suggests fascinating future directions for understanding uncertainty about the hypotheses in GVI given finite observations. The work in [6, 17] is concerned with small perturbations in the prior within a restricted class of measures $\mathcal{G}$. Often, restrictions are placed on this set in terms of how much the prior may vary; in [12] the priors considered are those close to the original prior $\Pi_{0}$ as $\varepsilon$ tends to zero. The present paper however is concerned with severe prior misspecification, even cases where we allow $\mathcal{G}=\mathcal{P}(\Theta)$, as well as the asymptotic dynamics of GVI posterior measures under such misspecification.
Appendix C Non–Robustness in Bayesian Inference
Continuing the illustrative example from Sections 2.3 and 4.2, with $X_{1}^{n}\overset{iid}{\sim}\mathcal{N}(\theta,\sigma_{p}^{2})$, we get the GVI objective. Using $\beta=1$ and loss $L(X_{1}^{n},\,\theta)=-\frac{1}{n}\log p_{\theta}(X_{1}^{n})$, $\mathbb{E}_{q(\theta)}[nL(X_{1}^{n},\,\theta)]$ is
$\mathbb{E}_{q(\theta)\sim\mathcal{N}(\mu_{q},\sigma_{q}^{2})}[-\log p_{\theta}(X_{1}^{n})]=\mathbb{E}_{q(\theta)}\left[-\log{(2\pi\sigma_{p}^{2})^{-\frac{n}{2}}}\exp\left\{-{\sum_{i=1}^{n}}\frac{(\theta-X_{i})^{2}}{2\sigma_{p}^{2}}\right\}\right]$
$\frac{n}{2}\log 2\pi\sigma_{p}^{2}+\frac{1}{2\sigma_{p}^{2}}\sum_{i=1}^{n}(\mu_{q}-X_{i})^{2}+\frac{n\sigma_{q}^{2}}{2\sigma_{p}^{2}}$
Therefore, we have a closed form for $R(D_{KL},-\frac{1}{n}\log p_{\theta},q)$ as
$\frac{n}{2}\log 2\pi\sigma_{p}^{2}+\frac{1}{2\sigma_{p}^{2}}\sum_{i=1}^{n}(\mu_{q}-X_{i})^{2}+\frac{n\sigma_{q}^{2}}{2\sigma_{p}^{2}}+\log\frac{\sigma_{\pi}}{\sigma_{q}}+\frac{(\mu_{q}-\mu_{\pi})^{2}+\sigma_{q}^{2}}{2\sigma_{\pi}^{2}}-\frac{1}{2}$
which is convex in both $\mu_{q}$ and $\sigma_{q}$. The minimisers of this objective are hence
$\nabla_{\mu_{q}}R(D_{KL},-\frac{1}{n}\log p_{\theta},q)$ $=\frac{1}{\sigma_{p}^{2}}\sum_{i=1}^{n}(\mu_{q}-X_{i})+\frac{\mu_{q}-\mu_{\pi}}{\sigma_{\pi}^{2}}=0$
$\nabla_{\sigma_{q}}R(D_{KL},-\frac{1}{n}\log p_{\theta},q)$ $=\frac{n\sigma_{q}}{\sigma_{p}^{2}}-\frac{1}{\sigma_{q}}+\frac{\sigma_{q}}{\sigma_{\pi}^{2}}=0$
Solving for $\sigma_{q}$ and $\mu_{q}$ gives the well known solution
$\sigma_{q}^{2}=\left(\sigma_{\pi}^{-2}+n\sigma_{p}^{-2}\right)^{-1},\,$ $\mu_{q}=\cfrac{\left(\frac{\mu_{\pi}}{\sigma_{\pi}^{2}}+\sum_{i=1}^{n}\frac{X_{i}}{\sigma_{p}^{2}}\right)}{\left(\sigma_{\pi}^{-2}+n\sigma_{p}^{-2}\right)}$
which shows that, for a finite $\sigma_{\pi}^{2}$, $\mu_{\pi}$ by itself can cause arbitrary misspecification.
{acks}
[Acknowledgments] The first author would like to direct correspondence to either Terje.Mildner@kellogg.ox.ac.uk or Terje.Mildner@gmail.com. {funding} TM and TD were supported by a UKRI Turing AI Acceleration Fellowship [EP/V02678X/1] and a Turing Impact Award from The Alan Turing Institute.

### v2

@@ 6.6. Rate of Convergence
${M}{\beta}^{-1}n^{-1}\geq Q_{n}N_{\varepsilon_{n}}^{c}(L_{n}(N_{\varepsilon_{n}}^{c})-L_{n}(\Theta))\geq 0\,.$
@@
6.7. Dependence on $n$
@@
This is immediate from the preceding proofs of Theorems 2 and 3. ∎
7. Concluding Remarks
@@
It is worth noting that these results have further impact to Federated GVI [mildner2025], the study of GVI posteriors over a network of clients with private, local data sets. In particular such FedGVI posteriors will exist under the assumptions presented here through the different choices of loss functions and priors. Moreover, a global prior of some uniformed server, given a bounded divergence, may not skew the global posterior after incorporating local data towards unfavourable hypotheses that do not agree.
Finally, while we were able to derive existence and uniqueness, as well as a consistency result for GVI posteriors under unbounded divergences, rates of convergence of GVI posterior measures under unbounded divergences remains an open challenge.
Acknowledgement
TM and TD were supported by a UKRI Turing AI Acceleration Fellowship [EP/V02678X/1] and a Turing Impact Award from The Alan Turing Institute. The first author would like to direct correspondence to Terje.Mildner@kellogg.ox.ac.uk or Terje.Mildner@gmail.com.
Appendix A General
@@
where the first is a Bochner integral.
@@ Appendix B Classical Robustness to Prior Misspecification
We briefly discuss the relation of our work to robustness to prior misspecification in the sense of Bayesian sensitivity analysis [berger1985, gustafson1995]. This line of work has examined the stability of Bayesian posteriors under small perturbations of the prior with respect to some other measure, for instance through additive Huber contamination as mentioned in Section 2.3. In particular, this has been used to select a robust prior for Bayes which leads to optimal decision making under potential perturbations of the prior measure. In this paper however, we aim to answer the reverse: under what conditions can GVI lead to optimal decision making under ill chosen priors.
Local sensitivity of posteriors generated with $\varepsilon$–contaminated priors, for instance by the additive Huber contamination model $\mathcal{G}=\{\Pi_{\varepsilon}:=(1-\varepsilon)\Pi_{0}+\varepsilon G:\varepsilon\in(0,1)\}$, is often studied [gustafson1995] through the Gâteux differential
@@
where $d(\cdot:\cdot)$ is some distance or divergence measure; Gustafson and Wasserman [gustafson1995] for instance use the total variation distance. Here $\Pi_{0}$ is the uncontaminated prior, $\Pi_{\varepsilon}$ the $\varepsilon$–contaminated prior in the direction of $G$, and $Q^{n}_{0}$ and $Q^{n}_{\varepsilon}$ the respective Bayes posteriors. The unrestricted choice of a divergence other than the Kullback–Leibler in GVI generally prohibits explicit expressions of $Q^{n}_{0}$ and $Q^{n}_{\varepsilon}$ that demonstrate the the exact dependence on the prior, which makes this approach impracticable for our purposes.
This presents an interesting theoretical question to the stability of Bayesian posteriors with respect to prior misspecification and suggests fascinating future directions for understanding uncertainty about the hypotheses in GVI given finite observations. The work in [berger1985, gustafson1995] is concerned with small perturbations in the prior within a restricted class of measures $\mathcal{G}$. Often, restrictions are placed on this set in terms of how much the prior may vary; in [dey1994] the priors considered are those close to the original prior $\Pi_{0}$ as $\varepsilon$ tends to zero. The present paper however is concerned with severe prior misspecification, even cases where we allow $\mathcal{G}=\mathcal{P}(\Theta)$, as well as the asymptotic dynamics of GVI posterior measures under such misspecification.